What is the probability that the mean random sample of 9 students will be less than 88? Given: = 4 ̅ = 88 = 86 n=9


1
Expert's answer
2022-06-16T06:29:11-0400
"P(X<88)=P(Z<\dfrac{88-\mu}{\sigma/\sqrt{n}})"

"=P(Z<\dfrac{88-86}{4/\sqrt{9}})=P(Z<1.5)"

"\approx0.9332"